OpenStudy (anonymous):

how many sides does a regular polygon have if the measure of one interior angle is 144

OpenStudy (anonymous):

10

OpenStudy (anonymous):

thanks do you have the work please ?

OpenStudy (anonymous):

okay so there's this formula to help you solve for the measurement of angles in equilateral polygons it goes like this; (n-2)*180 n stands for the number of sides so you could write it like this before solving it (n-2)*180=144

OpenStudy (anonymous):

\[\frac{ 180(n-2) }{ n }=144.\] Solve for n

OpenStudy (anonymous):

then solve for n

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

You are welcome

OpenStudy (anonymous):

no problem!

OpenStudy (vishweshshrimali5):

Well you can also derive the formula tayromo used. Try it and you will get a great deal of improvement in geometry

OpenStudy (anonymous):

can you slove it for me im still lost my teacher never thought us anything and finals are coming up :(

OpenStudy (vishweshshrimali5):

Its no problem

OpenStudy (vishweshshrimali5):

See the following figure:

OpenStudy (anonymous):

yes

OpenStudy (vishweshshrimali5):

Now here is one way to approach this problem: (1) You have to remember this property: Sum of exterior angles in a polygon is always equal to 360 degree

OpenStudy (vishweshshrimali5):

Do you remember this one ?

OpenStudy (anonymous):

yes

OpenStudy (vishweshshrimali5):

Gud

OpenStudy (vishweshshrimali5):

Now, let every interior angle be say x and number of sides be n. Now, since it is a regular polygon so every interioir angle will be equal. Right ?

OpenStudy (anonymous):